OpenStudy (pphalke):

If a right triangle has legs measuring 6 feet and 8 feet, how long is the hypotenuse?

OpenStudy (yoloshroom):

a^2 + b^2 = c^2 a and b are the legs, c is the hypotenuse 6^2 + 8^2 = c^2 100 = c^2 10 = c Im a bit rusty with math, but im about 90% sure that's how you do it.